OpenStudy (anonymous):

how to factor x^3+8????????????

OpenStudy (anonymous):

just like the last one

OpenStudy (anonymous):

\(a^3+b^3=(a+b)(a^2-ab+b^2)\) your job is to figure out what \(a\) and \(b\) are in this example it is pretty clear that \(a=x\) since \(a^3=x^3\) what is \(b\) ?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

oh no, \(b^3=8\) you need \(b\)

OpenStudy (anonymous):

\(x^3+8\) \[a^3+b^3\]

OpenStudy (anonymous):

b=2

OpenStudy (anonymous):

got it!

OpenStudy (anonymous):

so take \((a+b)(a^2-ab+b^2)\) and replace \(a\) by \(x\) and \(b\) by \(2\)

OpenStudy (anonymous):

(x+2)(x^2+2x+4)

OpenStudy (anonymous):

sorry (x+2)(x^2-2x+4)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is your picture a visual pun?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh, thought it might be "lulu's back, in town"

OpenStudy (anonymous):

haha i never thought of it that way, but i like it!